OpenStudy (anonymous):

Choose the prabability that cannot happen. A. 0.001 B. 1/16 C. 0.82 D. 5/4

jimthompson5910 (jim_thompson5910):

all probabilities are between 0 and 1 (inclusive)

jimthompson5910 (jim_thompson5910):

if p is some probability, then \[\Large 0 \le p \le 1\]

OpenStudy (anonymous):

Thank you it was d :)

jimthompson5910 (jim_thompson5910):

good, since 5/4 = 1.25 that is NOT in the interval from 0 to 1